OpenStudy (idek_73):

Given the following functions f(x) and g(x), solve (f + g)(3) and select the correct answer below: (1 point) f(x) = 2x + 21 g(x) = x − 24 −27 −21 6 48

Directrix (directrix):

(f + g) (3) = f(3) + g(3) = 2*3 + 12 + 3 - 24 = ? @idek_73

Directrix (directrix):

Post what you get, okay. When you post your next math question, please do so in the Math Section at http://openstudy.com/study#/groups/Mathematics The Feedback Forum (where we are now) is supposed to be for questions about the operation of OpenStudy. Welcome to OpenStudy i think you'll like the site.